# ViWikiFC: Fact-Checking for Vietnamese Wikipedia-Based Textual Knowledge Source

### v1

@@ 3.6.1 Overall Statistics
We divide the corpus, which consists of 20,916 claim sentences, randomly into three different sets: 80.00% for training (Train), 10.00% for development (Dev), and 10.00% for the test (Test) set for VP task. For ER tasks, we do not employ pre-trained models on a portion of the data and then evaluate the remaining part. Instead, we use the cosine similarity measure between sentences within the model to extract evidence sentences, and this process will be carried out and evaluated on the entire corpus. The distribution of labels is summarized in Table 4.
@@ 3.6.2 Length Distribution
Figure 4: The distribution total syllable in the corpus.
Figure 5: The distribution total word in the corpus.
3.6.3 New Word Rate
@@
[…]
Table 5: Corpus analysis in terms of linguistic aspects.
To evaluate diversity in the corpus, we measure the rate of new words in the claim sentence that do not appear in the evidence sentence. We use VnCoreNLP [56] for word segmentation. The results in table 5 show that the new word rate in the REFUTES label is the lowest, reaching only 31.28%. The word diversity of the NEI label (50.44%) is the highest among the three labels, followed by the SUPPORTS label with 32.91%. Both the SUPPORTS and REFUTES labels have a low new word rate compared to the NEI label, indicating that annotators tend to use fewer new words when constructing sentences for these two labels. This tendency ensures semantic accuracy with the evidence sentence when external knowledge is not used. In Addition, we utilized part-of-speech analysis on new words to delve deeper into annotator data construction trends. We used PhoNLP [57] for word segmentation and summarized the results in Table 5. The statistical results reveal that nouns and verbs are the two primary components annotators use when composing claim sentences.
3.6.4 New Dependency Rate
Figure 6: Semantic dependency tree from evidence to claim.
We measured the new dependency rate in all three labels to analyze the creativity in constructing claim sentences from evidence sentences. We use the VnCoreNLP toolkit [56] for dependency parsing on the claims and the evidence sentences. Subsequently, we calculated the dependencies in the claim sentences that were absent in the corresponding evidence sentences.
Figure 6 illustrates the modification of the semantic dependency tree from evidence to claim. The word in the red box appears in two sentences, and its position and dependency change throughout the syntax tree.
The results are summarized in Table 5. All three labels exhibit a new dependency rate of more than 50.00%, with NEI having the highest rate at 81.96%, followed by SUPPORTS at 75.26%, and REFUTES at 67.90%. The statistical results indicate that in the ViWikiFC dataset, the claim sentences constructed from the evidence sentences demonstrate the grammatical structure and semantic diversity. This contributes to making the ViWikiFC dataset more challenging when evaluating language models.
3.6.5 New n-gram rate
New n-gram rate ViWikiFC VitaminC
SUPPORTS REFUTES NEI SUPPORTS REFUTES NEI
Unigram 31.71 30.43 51.95 38.85 49.77 58.56
@@ 3.6.5 New n-gram rate
Table 6: New n-gram rate (%) between claim-evidence pairs.
Figure 7: The ratio of new n-grams number in ViWikiFC and VitaminC corpus.
To assess the similarity between the claim and evidence, we calculated the percentage of new n-grams between the claim and evidence sentences for different labels in our dataset. Furthermore, we have implemented a comparable approach to the VitaminC corpus [8] due to the resemblance between our corpus and the challenging corpus VitaminC. We used the nltk library  [^10] https://www.nltk.org/index.html to segment sentences into adjacent sequences of n syllables, ranging from one to four. The results are summarized in Table 6
From Table 6, across both corpus, the percentage of new n-grams is highest for the NEI label among the three labels. This is understandable because NEI sentences are written by annotators using additional information beyond the evidence sentence. In contrast, sentences with the SUPPORTS and REFUTES labels are written solely based on the information provided in the evidence, leading to a higher overlap of words or n-grams between the claim and evidence sentences, resulting in lower percentages of new n-grams in these two labels compared to the NEI label.
Furthermore, we also calculated the percentage distribution of the number of new n-grams in both data sets, as shown in Figure 7. The statistical results indicate that the distribution of new n-grams in the ViWikiFC data set is more evenly distributed. Sentences with more than 11 new n-grams have a high proportion in the dataset, in contrast to the VitaminC dataset, where the percentage is relatively small, indicating a more balanced distribution in the ViWikiFC dataset compared to VitaminC.
3.6.6 Corpus-generation rules analysis
To evaluate and analyze the linguistic factors of the annotators during the creation of the corpus, we proceed with the analysis of the rules used based on the ViNLI (Huynh et al.) corpus generation process. We randomly chose 500 SUPPORTS and 500 REFUTES claim-evidence pairs from the validation set. Selecting 500 sentences for each label will increase the coverage of the data compared to analyzing a smaller sample of only 100 sentences, as was performed in ViNLI.
No. Rule Ratio (%)
1 Converting active sentences to passive or vice versa 17.40
2 Replacing with synonyms 58.60
3 Adding/Removing modifier without changing the meaning of the evidence sentence 48.00
4 Turning nouns into relative clauses 10.20
5 Turning the object into relative clauses 2.20
6 Turning adjectives into relative clauses 0.20
7 Replacing quantifiers or time with others that have a similar meaning 5.60
8 Creating a presupposition sentence 20.60
9 Replacing Named Entities with a word that stands for the class 2.00
10 Creating conditional sentences 0.40
11 Others 3.00
Table 7: Data-generation rules for creating SUPPORTS claims based on evidence.
No. Rule Ratio (%
1 Using negative words (cannot, hardly, do not,…) 23.00
2 Replacing with antonyms 14.00
3 Opposite of quantity 9.80
4 Opposite of time 5.60
5 Creating a sentence that has the opposite meaning of a presupposition 5.60
6 Wrong reasoning about an object (House, car, river, sea, person, etc.) 23.60
7 Wrong reasoning about an event 30.80
8 Others 0.20
Table 8: Data-generation rules for creating REFUTED claims based on evidence.
For the SUPPORTS label, the rule annotator uses most is ’replace words with synonyms’ with 58.60%; meanwhile, the ’turn adjective into relative clause’ rule has the lowest ratio (Table 7). As for the REFUTES label, annotators tend to generate claims with the ’wrong reasoning about an event’ rule, whereas ’opposite of time’ and ’create a sentence that has the opposite meaning of a presupposition’ share the same proportion of 5.60%, making them the least common rule used when creating data (Table 8).
Figure 8: The ratio of SUPPORT claim generation rules.
Figure 9: The ratio of REFUTE claim generation rules.
During the corpus generation process, annotators can use more than one rule for every claim they write. The analysis we conducted on claims shows that almost half of SUPPORT claims are written using two rules (49.80%), while one rule has 41.20%. For the REFUTES label, claims are nearly generated by one rule with a proportion of 88.00% (Figure 9). This can be explained by the fact that when using too many rules without reviewing carefully, REFUTES claims can be turned into NEI labels (Table 3).
4 Empirical Evaluation
4.1 Baseline Models
To assess the challenge of ViWikiFC, we define two tasks for a comprehensive corpus evaluation: evidence retrieval and verdict prediction (VP). Moreover, we integrate these two tasks into a complete pipeline. Each task is evaluated on the development set, and the final prediction results are based on the test set.
4.1.1 Evidence Retrieval
For the evidence retrieval (ER) task, we give a claim sentence as input, then employ TF-IDF, Okapi BM25, and Vietnamese Sentence-BERT [10] to select the top one to ten sentences with the highest relevance to the claim sentence from the extracted corpus. With TF-IDF, we build the algorithm from scratch, while with BM25, we use the Rank-BM25111111https://pypi.org/project/rank-bm25/ library with pre-set hyperparameters.
Vietnamese Sentence-BERT [10] is a Vietnamese Sentence embedding model using sentence-BERT [58] architecture. Sentence-BERT [58] is a variant of the pre-trained model BERT [13] with a siamese and triplet network structure, which enhances the ability to search and compare the correlation between sentences with reduced computational time and cost compared to other pre-trained models such as BERT [13] and RoBERTa [49].
4.1.2 Verdict Prediction
In the VP task, we conducted experiments using models ranging from simple approaches like Continuous Bag of Words (CBOW) to more complex ones like BiLSTM [12]. These two models are used with the pre-trained PhoW2V embedding [59] for Vietnamese. We applied two 300-dimensional versions of PhoW2V, including the syllable and word levels. Furthermore, we utilize state-of-the-art multilingual pre-trained language models such as Multilingual BERT (mBERT) [13], XLM-RoBERTa [14], InfoXLM [15], as well as monolingual pre-trained language models for Vietnamese such as PhoBERT [16], ViT5 [17], and ViDeBERTa [18].
The six pre-trained models, mBERT, XLM-R, PhoBERT, InfoXLM, ViT5, and ViDeBERTa, perform highly on various natural language processing (NLP) tasks, including verdict prediction. mBERT, XLM-R, and InfoXLM are multilingual models pre-trained on a vast amount of data encompassing over 90 languages, including Vietnamese. PhoBERT, on the other hand, is a state-of-the-art pre-trained language model for Vietnamese, trained on 20GB Vietnamese Wikipedia and news text corpus, with 135 million parameters for the base version, 370 million parameters for the large version, and recently, PhoBERT base-v2 version have been published which trained on 120GB of Vietnamese texts from OSCAR-2301121212https://huggingface.co/datasets/oscar-corpus/OSCAR-2301. ViT5 follows encode-decoder transformer architecture [48] with T5 [51] self-supervised pertaining framework. ViDeBERTa is pre-trained with DeBERTa architecture [52] trained on a 138GB Vietnamese dataset.
A characteristic of the Vietnamese language is that a word can be composed of various syllables, so to evaluate various aspects of the Vietnamese language in this task, we conducted two experiments using word-based (PhoBERT, ViDeBERTa) and syllable-based (mBERT, XLM-R, InfoXLM, ViT5) inputs for the models. Word-based models are used with the VNCoreNLP toolkit [56] for word segmentation. Furthermore, to mitigate the influence of the experiment results that vary due to initialization seeds, we conducted the experiments five times with randomly selected seeds. We calculated the standard deviation of the results obtained.
4.1.3 Pipeline Description
Finally, we selected models that achieved the highest performance in both tasks, evidence retrieval and verdict prediction. To ensure the best rigor and accuracy for the pipeline, we use strict accuracy measurement where the input is a sentence (claim), and the output consists of a sentence (evidence) extracted from the corpus along with a label of the claim validated based on information within the evidence sentence. The pipeline will count as incorrect output whenever the evidence or label is incorrect. Strict accuracy is described below, with v and v’ as the ground truth and predicted evidence in the evidence retrieval task, e and e’ as the ground truth and predicted label in verdict prediction (e, e’ $\in$ SUPPROTS, REFUTES, NOTENOUGHINFORMATION)
Strict Accuracy = $\delta(v,v^{\prime})\times\delta(e,e^{\prime})$
4.1.4 Evaluation Metrics
F1 = $\frac{2\ \times\ Precision\ \times\ Recall}{Precision\ +\ Recall}$
4.2 Experimental Results
4.2.1 Evidence Retrieval
@@
We can observe that BM25 consistently outperforms TF-IDF and SBERT across all 3 labels in both single-sentence and 5-sentence retrieval scenarios. As the number of retrieved sentences increases from 1 to 5, the accuracy of all 3 models increases, with BM25 achieving the highest accuracy, reaching 93.93% for the SUPPORTS label and 93.04% for the REFUTES label. SBERT, with its transformer architecture, possesses the ability to capture the contextual nuances of the sentences it retrieves, but in the ViWikiFC corpus, the model’s performance underperforms BM25. This suggests that BM25, despite its less complex architecture, exhibits remarkable performance in evidence retrieval tasks compared to other transformer-based methods [60].
@@ 4.2.2 Verdict Prediction
In the case of the ”word” corpus, PhoBERTLarge stands out as the best-performing model. It achieves an impressive accuracy of 82.88% on the development set and 81.63% on the test set. These results demonstrate the robustness and effectiveness of PhoBERTLarge in handling the nuances of the ”word” corpus. Its high accuracy in both development and test sets suggests that it can consistently provide reliable predictions in this specific context.
However, when we shift our focus to the ”syllable” corpus, InfoXLMLarge takes the lead with remarkable accuracy and F1 scores. It achieves an accuracy of 87.45% and 86.50% on validation and test sets, respectively. These numbers not only surpass the performance of PhoBERTLarge but also highlight the exceptional capabilities of InfoXLMLarge in handling the ”syllable” corpus. This outcome establishes InfoXLMLarge as the optimal model for the VP task. The superiority of InfoXLMLarge in the ”syllable” corpus might be attributed to its ability to capture the intricate linguistic patterns and syllable-level features, which are evidently critical in achieving high accuracy and F1 scores in this particular context.
Model Dev Test
Acc F1 Acc F1
Word CBoW 51.26±$\pm{0.66}$ 50.50±$\pm{0.98}$ 50.39±$\pm{1.05}$ 49.65±$\pm{0.95}$
[…]
Syllable CBoW 45.25 ±$\pm{0.66}$ 44.55 ±$\pm{0.63}$ 44.24 ±$\pm{1.05}$ 43.46 ±$\pm{0.64}$
[…]
Table 10: The result of models on Dev and Test.
4.2.3 Pipeline
@@
Table 11 gives the result of the pipeline; the best combination for both tasks is BM25 and InfoXLMLarge with 67.00%, while SBERT and PhoBERTLarge have the lowest result. This result demonstrates ViWikiFC corpus’s challenge for current Vietnamese fact-checking models. Besides, it provides a practical perspective on the fact-checking problem, as accurate evidence retrieval significantly influences the performance of evidence-based methods such as PhoBERT and InfoXLM. While these methods perform well in verdict prediction tasks, their strict accuracy still needs to be improved. Therefore, approaching the Vietnamese fact-checking problem using separate models for each task, first for evidence retrieval and then for verdict prediction, needs improvement to deals with more complex claim that have been studied in English [4, 61].
4.3 Result Analysis
4.3.1 Evidence Retrieval
Figure 10: The effect of new POS tagging for ER task.
Figure 10 illustrates how model accuracy changes in the ER task as the number of new words increases. The results consistently show a decline in models performance as the count of unfamiliar words in various word categories rises. In particular, we can observe in the chart depicting the number of new preposition that the BM25 model significantly decreases from more than 70.00% to approximate 0% as we increase the number of preposition.
When we introduce a larger number of new words, especially a preposition, there can be significant changes in the context of the sentence. This can pose challenges to both BM25 and SBERT model in accurately determining the context for the information that needs to be retrieved, potentially leading to the return of unrelated or inaccurate sentences and reducing the accuracy of the model.
Furthermore, most sentences that contain many new words belong to the NEI label (see Table 5). During the process of generating NEI-labeled claims, annotators use information and details that are not related to the evidence, making it challenging for the model to retrieve within the scope of a single sentence accurately.
4.3.2 Verdict Prediction
To gain a clearer understanding of the data in the verdict prediction task and approach the language from different aspects, we utilize the two best models, PhoBERTLarge and InfoXLMLarge. In this section, InfoXLM and PhoBERT respectively replace PhoBERTLarge and InfoXLMLarge. We manually determined the number of rules employed to create each sentence (as mentioned in section 3.6.6).
Label Rule PhoBERT InfoXLM
SUPPORTS 1 79.25 88.35
[…]
REFUTES 1 80.68 82.95
[…]
Table 12: Impacts of data-generation rules on the models.
Figure 11: The effect of claim length
Figure 12: The effect of new word rate for each label in verdict prediction task
As we can see in Table 4.3.2, SUPPORT claims that are constructed using only one rule can indeed pose challenges for the model when predicting the correct label. As for the REFUTES label, similar to the SUPPORTS label, the predictions are divergent between the two models when applied to two types of different label quantities. For PhoBERT, making substantial errors in information tends to make it easier for models to detect discrepancies compared to cases where they need to identify just one incorrect piece of information between the claim and the original evidence.
We also aim to gain insights into how sentence length affects the prediction results of the model. Figure 11 illustrates how the accuracy of the models changes as we increase the length of the sentences. The accuracy of the model increases as the sentences become longer. It is apparent that there is a positive correlation between sentence length and model accuracy. The trend suggests that the model benefits from having more context to make accurate predictions. Longer sentences provide additional context and information, allowing the model to understand the context of the text better and, consequently, make more precise predictions.
In general, for both the SUPPORTS and REFUTES labels, the accuracy of the models tends to decrease as the percentage of new words increases (see Figure 12). Surprisingly, in the SUPPORTS label, when the percentage of new words is in the range of 41.00-60.00%, InfoXLM yields slightly better results compared to the previous range of 21.00-40.00%. PhoBERT has significantly lower accuracy than InfoXLM in correctly predicting the SUPPORTS label in the range of 4.00-13.00%. The REFUTES label shows a clear downward trend with InfoXLM as the percentage of new words in the claim sentence increases. However, as the percentage of 41.00-60.00% new words increases, it becomes evident that PhoBERT predicts the label slightly better than InfoXLM. The NEI label is characterized by a high percentage of new words, so it is not surprising that as the number of new words increases, the model’s prediction accuracy also increases significantly. In terms of accuracy, we can observe that InfoXLM performs better than PhoBERT for percentages of new words below 60.00%. Afterward, both models exhibit similar performance in the remaining portion of the dataset.
5 Conclusions and Future Directions

### v2

@@ 3.5.1.  Overall Statistics
We divide the corpus, which consists of 20,916 claim sentences, randomly into three different sets: 80.00% for training (Train), 10.00% for development (Dev), and 10.00% for the test (Test) set for the VP task. For ER tasks, we do not employ pre-trained models on a portion of the data and then evaluate the remaining part. Instead, we use the cosine similarity measure between sentences within the model to extract evidence sentences, and this process will be carried out and evaluated on the entire corpus. The distribution of labels is summarized in Table 2.
@@ 3.5.2.  Length Distribution
Figure 2: The distribution of total syllables in the corpus.
Figure 3: The distribution of total words in the corpus.
3.5.3.  New Word Rate
@@
REFUTES 31.28 67.90 26.27 25.47 8.66 8.04 13.10 18.47
NEI 50.44 81.96 33.96 23.61 8.96 9.82 7.65 15.99
Table 3: Corpus analysis in terms of linguistic aspects.
To evaluate diversity in the corpus, we measure the rate of new words in the claim sentence that do not appear in the evidence sentence. We use VnCoreNLP Vu et al. (2018) for word segmentation. The results in table 3 show that the new word rate in the REFUTES label is the lowest, reaching only 31.28%. The word diversity of the NEI label (50.44%) is the highest among the three labels, followed by the SUPPORTS label with 32.91%. Both the SUPPORTS and REFUTES labels have a low new word rate compared to the NEI label, indicating that annotators tend to use fewer new words when constructing sentences for these two labels. This tendency ensures semantic accuracy with the evidence sentence when external knowledge is not used. In Addition, we utilized part-of-speech analysis on new words to delve deeper into annotator data construction trends. We used PhoNLP Nguyen and Nguyen (2021) for word segmentation and summarized the results in Table 3. The statistical results reveal that nouns and verbs are the two primary components annotators use when composing claim sentences.
3.5.4.  Corpus-generation rules analysis
To evaluate and analyze the linguistic factors of the annotators during the creation of the corpus, we proceed with the analysis of the rules used based on the ViNLI Huynh et al. (2022) corpus generation process. We randomly chose 500 SUPPORTS and 500 REFUTES claim-evidence pairs from the validation set. Selecting 500 sentences for each label will increase the coverage of the data compared to analyzing a smaller sample of only 100 sentences, as was performed in ViNLI. The statistical results and illustrative examples are provided in Appendix A.
For the SUPPORTS label, the rule annotator uses most is ’replace words with synonyms’ with 58.60%; meanwhile, the ’turn adjective into relative clause’ rule has the lowest ratio (Table LABEL:tab:_support_rules_and_example). As for the REFUTES label, annotators tend to generate claims with the ’wrong reasoning about an event’ rule, whereas ’opposite of time’ and ’create a sentence that has the opposite meaning of a presupposition’ share the same proportion of 5.60%, making them the least common rule used when creating data (Table LABEL:tab:_refute_rules_and_example).
Figure 4: The ratio of SUPPORT claim generation rules.
Figure 5: The ratio of REFUTE claim generation rules.
During the corpus generation process, annotators can use more than one rule for every claim they write. The analysis we conducted on claims shows that almost half of SUPPORT claims are written using two rules (49.80%), while one rule has 41.20%. For the REFUTES label, claims are nearly generated by one rule with a proportion of 88.00% (Figure 5). This can be explained by the fact that when using too many rules without reviewing carefully, REFUTES claims can be turned into NEI labels.
4.  Empirical Evaluation
4.1.  Baseline Models and Pipeline Description
To comprehensively evaluate the ViWikiFC dataset, we define two main subtasks: Evidence Retrieval (ER) and Verdict Prediction (VP), which are further integrated into a complete fact-checking pipeline.
4.1.1.  Evidence Retrieval
In the ER task, given a claim sentence, the objective is to identify the most relevant evidence sentences from the corpus. We employ three retrieval methods: TF-IDF, Okapi BM25, and Vietnamese Sentence-BERT (SBERT). The TF-IDF model is implemented from scratch, BM25 is applied using the Rank-BM25555https://pypi.org/project/rank-bm25/ library with default hyperparameters, and Vietnamese Sentence-BERT is adopted to capture semantic similarity between claims and evidence through sentence embeddings. This model, based on the Sentence-BERT architecture, enables efficient comparison of sentence-level meanings with reduced computational cost compared to traditional transformer-based encoders.
4.1.2.  Verdict Prediction
For the VP task, we explore a range of models from traditional neural architectures to pre-trained language models. Simpler models, such as Continuous Bag of Words (CBOW) and BiLSTM, are trained with pre-trained PhoW2V embeddings at both syllable and word levels. We further investigate the performance of advanced transformer-based models, including Multilingual BERT (mBERT), XLM-RoBERTa, InfoXLM, PhoBERT, and ViDeBERTa, covering both multilingual and monolingual representations for Vietnamese. Word-based models (PhoBERT, ViDeBERTa) employ the VNCoreNLP toolkit for word segmentation, while syllable-based models (mBERT, XLM-R, InfoXLM) use raw tokenized text. To ensure robustness, each experiment is conducted five times with different random seeds, and the standard deviation of the performance metrics is reported.
4.1.3.  Pipeline Description
Finally, we selected the models that achieved the highest performance in both tasks, evidence retrieval and verdict prediction, and integrated them into a complete fact-checking pipeline. This pipeline takes a claim sentence as input and outputs both an evidence sentence extracted from the corpus and a corresponding verdict label that indicates whether the claim is SUPPORTED, REFUTED, or NOT ENOUGH INFORMATION.
To ensure rigorous and reliable evaluation, we apply a strict accuracy metric, which considers a prediction correct only when both the retrieved evidence and the predicted label are accurate. In other words, if either the evidence or the verdict is incorrect, the entire prediction is counted as incorrect. This metric reflects the end-to-end quality of the fact-checking process rather than evaluating each component in isolation.
Strict accuracy is formally defined as follows, where $v$ and $v^{\prime}$ denote the ground-truth and predicted evidence in the evidence retrieval task, and $e$ and $e^{\prime}$ represent the ground-truth and predicted labels in the verdict prediction task $e,e^{\prime}\in\{SUPPORTS,REFUTES,NOTENOUGHINFORM\\ ATION\}$:
Strict Accuracy = $\delta(v,v^{\prime})\times\delta(e,e^{\prime})$
where $\delta(a,b)$ equals 1 if $a=b$ and 0 otherwise.
4.1.4.  Evaluation Metrics
F1 = $\frac{2\ \times\ Precision\ \times\ Recall}{Precision\ +\ Recall}$
4.2.  Experimental Results
Model Dev Test
Acc F1 Acc F1
@@ 4.2.  Experimental Results
BiLSTM 49.91 $\pm{0.47}$ 49.83 $\pm{0.61}$ 49.13 $\pm{0.45}$ 49.24 $\pm{0.48}$
@@
PhoBERTBaseV2 82.50 $\pm{0.52}$ 82.28 $\pm{0.53}$ 81.22 $\pm{0.53}$ 81.19 $\pm{0.56}$
PhoBERTLarge 82.88 $\pm{0.97}$ 82.74 $\pm{1.01}$ 81.63 $\pm{0.47}$ 81.62 $\pm{0.44}$
Syllable CBoW 45.25 $\pm{0.66}$ 44.55 $\pm{0.63}$ 44.24 $\pm{1.05}$ 43.46 $\pm{0.64}$
BiLSTM 47.68 $\pm{0.60}$ 47.67 $\pm{0.77}$ 45.12 $\pm{0.62}$ 45.04 $\pm{0.79}$
mBERT 75.75 $\pm{0.68}$ 75.61 $\pm{1.24}$ 75.93 $\pm{0.58}$ 76.01 $\pm{1.18}$
XLM-RBase 79.66 $\pm{0.95}$ 79.48 $\pm{0.97}$ 78.53 $\pm{0.41}$ 78.56 $\pm{0.48}$
XLM-RLarge 86.09 $\pm{0.72}$ 86.02 $\pm{0.67}$ 85.11 $\pm{0.27}$ 85.15 $\pm{0.26}$
InfoXLMBase 80.10 $\pm{0.57}$ 79.92 $\pm{0.61}$ 79.19 $\pm{0.78}$ 79.23 $\pm{0.77}$
InfoXLMLarge 87.45 $\pm{0.41}$ 87.42 $\pm{0.60}$ 86.50 $\pm{0.43}$ 86.51 $\pm{0.62}$
Table 4: The result of models on Dev and Test.
4.2.1.  Evidence Retrieval
@@
We can observe that BM25 consistently outperforms TF-IDF and SBERT across all 3 labels in both single-sentence and 5-sentence retrieval scenarios. As the number of retrieved sentences increases from 1 to 5, the accuracy of all 3 models increases, with BM25 achieving the highest accuracy, reaching 93.93% for the SUPPORTS label and 93.04% for the REFUTES label. SBERT, with its transformer architecture, possesses the ability to capture the contextual nuances of the sentences it retrieves, but in the ViWikiFC corpus, the model’s performance underperforms BM25. This suggests that BM25, despite its less complex architecture, exhibits remarkable performance in evidence retrieval tasks compared to other transformer-based methods Thakur et al. (2021).
@@ 4.2.2.  Verdict Prediction
In the case of the "word" corpus, PhoBERTLarge stands out as the best-performing model. It achieves an impressive accuracy of 82.88% on the development set and 81.63% on the test set. These results demonstrate the robustness and effectiveness of PhoBERTLarge in handling the nuances of the "word" corpus. Its high accuracy in both development and test sets suggests that it can consistently provide reliable predictions in this specific context.
However, when we shift our focus to the "syllable" corpus, InfoXLMLarge takes the lead with remarkable accuracy and F1 scores. It achieves an accuracy of 87.45% and 86.50% on validation and test sets, respectively. These numbers not only surpass the performance of PhoBERTLarge but also highlight the exceptional capabilities of InfoXLMLarge in handling the "syllable" corpus. This outcome establishes InfoXLMLarge as the optimal model for the VP task. The superiority of InfoXLMLarge in the "syllable" corpus might be attributed to its ability to capture the intricate linguistic patterns and syllable-level features, which are evidently critical in achieving high accuracy and F1 scores in this particular context.
4.2.3.  Pipeline
@@
Table 6 gives the result of the pipeline; the best combination for both tasks is BM25 and InfoXLMLarge with 67.00%, while SBERT and PhoBERTLarge have the lowest result. This result demonstrates the ViWikiFC corpus’s challenge for current Vietnamese fact-checking models. Besides, it provides a practical perspective on the fact-checking problem, as accurate evidence retrieval significantly influences the performance of evidence-based methods such as PhoBERT and InfoXLM. While these methods perform well in verdict prediction tasks, their strict accuracy still needs to be improved. Therefore, approaching the Vietnamese fact-checking problem using separate models for each task, first for evidence retrieval and then for verdict prediction, needs improvement to deal with more complex claims that have been studied in English Pan et al. (2023); Yao et al. (2023).
5.  Conclusions and Future Directions